Fermi National Laboratory
 

Operation Summary for October 5, 2009 - October 7, 2009

 

Monday Oct. 5

The day shift began with a Tevatron (TeV) quench investigation, Operations monitoring the Recycler stash (99.4E10), with the Antiproton Source stacking (11.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, TeV personnel continued the sector F13 quench investigation. Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Mechanical support personnel made an access into F-Sector to connect a vacuum pump at F13. Operations prepared the following areas for access: Pre-Acc, MiniBooNE, Transport, CDF, DZero, and the Pbar Rings. Efforts to restore beam to Pbar, NuMI, and MiniBooNE were underway by the end of the shift.

During the evening shift, an update to the TeV sector C refrigerator controls hardware became detrimental to sector C cryo system. Cryo recovery lasted until 9 PM. Operations established beam to NuMI at 4:45 PM. Pbar and MiniBooNE beam operations resumed by 5:15 PM.

Tuesday Oct. 6

The midnight shift began with Operations monitoring the stash (277.8E10), with Pbar stacking (26.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TeV into HEP shot setup at 12:15 AM and established store 7243 at 2:10 AM with an average initial luminosity (AIL) of 229.89E30. The B48 separator positive plate power supply tripped off. Operators adjusted the negative plate power supply output voltage to minimize losses in the TeV.

During the day shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Store 7243 collided until the scheduled termination time of 2:30 PM. HV repairs on the B48+ separator followed and were completed by 3:30 PM.

During the evening shift, Operations put the TeV into HEP shot setup at 5:00 PM and established store 7246 at 6:25 PM with an average initial luminosity (AIL) of 208.01E30.

Wednesday Oct. 7

The midnight shift began with Operations monitoring store 7246 (102.50) and the stash (152.5E10), with Pbar stacking (24.2E10), and with NuMI and MiniBooNE taking beam.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/07/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory